OpenStudy (anonymous):

x-4 square root x-2=2

OpenStudy (anonymous):

do you mean x=4?

OpenStudy (anonymous):

no, like x-4 then the square root of x-2=2

OpenStudy (anonymous):

as in \[x-(4\sqrt{x-2})=2\]

OpenStudy (radar):

\[(x-4)\sqrt{x-2}\] Is that the way it looks?

OpenStudy (anonymous):

yes but after the 2 their is an =2

OpenStudy (radar):

I can't work it, I end up with a 3rd degree polynomial !! Can't figure it out, it looks so simple. I feel sort of foolish about.

OpenStudy (anonymous):

if you're tying to solve for x then this gives an approximate answer (using computer programming) as 5.1304 however, I am more inclined to believe that the question contains a misprint because it is impossible to solve for x exactly.